OpenStudy (anonymous):

Determine the number of solutions of the linear system: 14x + 7y = 315 16x - 2y = 610

OpenStudy (anonymous):

elimination method or substitution

OpenStudy (anonymous):

Since the system is linear and the lines do not intersect, there can only be one solution.

OpenStudy (anonymous):

if they do not intersect then they have no solution...

OpenStudy (anonymous):

if you multiply the first by 2 and the second by 7 you get \[28x+14y=630\] \[112x-14y=4270\] adding these you get \[140x=4900\] \[x=35\]

OpenStudy (anonymous):

find y by plugging in x=35 to one of the above equations.. This makes total sense because both are linear equations, so they're lines. the lines will meet only at one point because if you continuously draw a two lines.

OpenStudy (anonymous):

they will only hit once, never hit, or be the same line

OpenStudy (anonymous):

does this make sense?

OpenStudy (anonymous):

well yes.. but I'm a little confused... is it possible to determine if they only hit once, never hit, or are the same line?

OpenStudy (anonymous):

yes if they're the same line, one line is a linear combination of the other. What this means is \[c_1x+c_2y=k\]

OpenStudy (anonymous):

so say you have a two lines \[2x+4y=10\] \[x+2y=5\] these are the same line because if you multiply the bottom by 2 you get \[2x+4y=10\] \[2x+4y=10\]

OpenStudy (anonymous):

in this case if you have two exact lines, they hit the same exact points infinitely which means there is tons of solutions

OpenStudy (anonymous):

No solution happens when you have two lines that when you add them using elimination you cancel one side out and you get a false statement. An example of this is two lines with the same slope having different y intercepts

OpenStudy (anonymous):

\[2y+2x=400\] \[2y+2x=500\] if you multiply the second by -1 and add them you get \[0=-100\] this is false because 0 does not equal -100. In this case there is no solution the picture of this would be 

OpenStudy (anonymous):

lastly like your question if when you eliminate you're able to solve for one variable, there is one solution to the system

OpenStudy (anonymous):

\[2x+4y=10\] \[x+5y=25\] multiply the last by -2 and add both equations \[2x+4y=10\] \[-2x-10y=-50\] \[-6y=-40\] \[y=40/6\]

OpenStudy (anonymous):

now use this in one of thebeginning equations to find x i'll pick the first one \[2x+4(40/6)=10\] \[2x+(160/6)=(60/6)\] \[2x=(-100/6)\] \[x=(-100/12)\]

OpenStudy (anonymous):

so the point at which they intersect is \[(-100/12,40/6)\]

OpenStudy (anonymous):

Okay great answer!! Thank you :)

OpenStudy (anonymous):

no proble

OpenStudy (anonymous):

could you also answer my other graphing question ?? haha.. :)